OpenStudy (anonymous):

solve the radical equation for x 2x - 10 square root the whole thing and then it = 8

OpenStudy (anonymous):

ANSWER IS 37

jimthompson5910 (jim_thompson5910):

sqrt(2x-10)=8 2x-10=8^2 2x-10=64 2x=64+10 2x=74 x=74/2 x=37

OpenStudy (anonymous):

IF ANSWER IS CORRECT, GIVE THE MEDALS